Solve the equation w=Cr^-4 for r
Answer by asuar010(338)

the first thing to do is to rewrite the equatoin as w=C/r^4...
then we multiply both sides by r^4 and we obtain r^4*w=C...
then we divide both sides by w and we obtain r^4=C/w...
the we take the fourth root of both sides and we obtain r=(C/w)^(1/4)